OpenStudy (anonymous):

square root of 512m^3

OpenStudy (swissgirl):

\(\sqrt{512*m*m*m}\) \(m\sqrt{16*32*m}\) \(4m\sqrt{8*4*m}\) \(2*4*m\sqrt{2*4*m}\) \2*2*4*m\sqrt{2m}\) \(16m\sqrt{2m}\)